# Evolutionary Policy Optimization

### v1

@@ Abstract
Despite its extreme sample inefficiency, on-policy reinforcement learning has become a fundamental tool in real-world applications. With recent advances in GPU-driven simulation, the ability to collect vast amounts of data for RL training has scaled exponentially. However, studies show that current on-policy methods, such as PPO, fail to fully leverage the benefits of parallelized environments, leading to performance saturation beyond a certain scale. In contrast, Evolutionary Algorithms (EAs) excel at increasing diversity through randomization, making them a natural complement to RL. However, existing EvoRL methods have struggled to gain widespread adoption due to their extreme sample inefficiency. To address these challenges, we introduce Evolutionary Policy Optimization (EPO), a novel policy gradient algorithm that combines the strengths of EA and policy gradients. We show that EPO significantly improves performance across diverse and challenging environments, demonstrating superior scalability with parallelized simulations. For visualizations of the learned policies, please visit: https://sites.google.com/view/epo-rl.
@@ 1 Introduction
Reinforcement Learning (RL) has emerged as a powerful framework for training autonomous decision-making agents in various domains, including games (Silver et al., 2016; Berner et al., 2019; Vinyals et al., 2019; Mnih, 2016), robotics (Kalashnikov et al., 2018; Akkaya et al., 2019; Kumar et al., 2021; Hwangbo et al., 2019), and large language models (LLMs) (Ouyang et al., 2022; Guo et al., 2025; Team et al., 2023). RL, whether off-policy or on-policy, follows a trial-and-error approach and is inherently sample inefficient. On-policy RL methods, in particular, suffer more from inefficiency compared to off-policy or model-based alternatives since they discard past experience.
Despite their sample inefficiency, on-policy model-free RL methods, also known as policy gradients(Schulman et al., 2017; Mnih, 2016), are widely adopted in real-world applications (Silver et al., 2016; Berner et al., 2019; Kalashnikov et al., 2018; Akkaya et al., 2019; Ouyang et al., 2022). Their popularity stems from their ability to achieve higher asymptotic rewards and stable performance, making them the preferred choice in data-rich environments such as simulations and game engines (Makoviychuk et al., ), where large training batches can be easily generated. However, recent studies(Singla et al., 2024; Petrenko et al., 2023) reveal a critical limitation: on-policy RL methods do not scale well with increasing batch sizes. Beyond a certain point, performance plateaus despite additional data collection, indicating that simply generating more samples across GPUs does not necessarily improve learning.
This limitation arises from the fundamental nature of on-policy RL: since data is collected from the current policy, increasing the number of parallel environments does not guarantee more diverse experiences. Instead, data distributions tend to converge, reducing the benefits of additional sampling. Evolutionary Algorithms (EAs), in contrast, excel at increasing diversity through randomization, making them a natural complement to RL. Evolutionary Reinforcement Learning (EvoRL), which integrates EAs with RL, is not a new concept (Moriarty et al., 1999) and has shown promise in prior work (Salimans et al., 2017). By maintaining a population of diverse policies and refining them through evolutionary operations like crossover and mutation, EvoRL enhances policy search and exploration.
However, EvoRL has not gained mainstream adoption due to its extreme sample inefficiency. In its simplest form, EvoRL resembles random search (Mania et al., 2018), as it lacks direct reward optimization like policy gradients. Typical EvoRL methods optimize reward indirectly through gradient-free search, known as Evolutionary Strategies (ES) (Khadka & Tumer, 2018; Chen et al., 2019; Zheng & Cheng, 2023), but training large deep models with ES is computationally prohibitive due to excessive memory and compute requirements. Even hybrid approaches that update a master agent via gradient ascent (Pourchot & Sigaud, 2019) struggle with efficiently maintaining and updating a large population of agents within GPU constraints, leading to low-quality data generation and poor sample efficiency. A notable hybrid approach, Population-Based Training (PBT) (Petrenko et al., 2023), aims to combine policy gradients with EvoRL by evolving a population of PPO policies. However, PBT segregates each policy in the population, preventing effective amortization of experience across agents and ultimately limiting its performance.
@@
In this paper, we introduce Evolutionary Policy Optimization (EPO), a novel policy gradient algorithm that combines the strengths of EA and policy gradients: 1) EA to increase the diversity of experience in massively parallelized training 2) Amortized experience from all policies in a population to directly train a master policy. To achieve this, we adopt the split and aggregate policy gradients (SAPG) formulation (Singla et al., 2024), which performs off-policy updates via importance sampling to aggregate experience from individual follower policies into the on-policy data of the master policy.
@@
EPO maintains a population of policies alongside a central master policy, all sharing the same network weights (analogous to gene-expression rules). This prevents parameter bloat as the population scales, enabling efficient training and storage of large networks. Each agent is conditioned on a unique latent variable (”gene”), ensuring behavioral diversity while maintaining coherence across policies. Periodically, Darwinian natural selection (Darwin, 2023) is applied: low-performing agents are eliminated, while top-performing elite agents are preserved. These elites undergo crossover and mutation, injecting controlled diversity while avoiding excessive divergence. All policies are updated synchronously via policy gradients, using shared reward signals to ensure efficient evolution, allowing EPO to scale effectively with larger neural networks. In each iteration, EPO aggregates experience from all policies into the master agent via SAPG updates (Singla et al., 2024), enabling the master policy to learn from diverse, high-quality data. This allows EPO to scale efficiently with increasing batch sizes, fully leveraging the potential of massively parallel simulations.
@@
We evaluate EPO on several challenging RL benchmarks and demonstrate that it significantly outperforms prior state-of-the-art RL methods, including PBT (Petrenko et al., 2023) and SAPG (Singla et al., 2024). Furthermore, we provide scaling laws, showing that EPO effectively scales with increased computational resources, making it a promising approach for large-scale RL training.
@@ 2 Related Work
On-policy approaches (Sutton et al., 1999) update the policy using data collected by the current policy, ensuring that the training data remains closely aligned with the policy’s behavior. This strategy often leads to higher asymptotic rewards and stable performance, but at the cost of sample inefficiency, as past experiences from older policies are discarded. Despite this drawback, on-policy methods are the preferred choice in settings where data is abundant, such as simulations and game engines, where large training batches can be efficiently generated. In practice, on-policy RL has proven highly effective for training autonomous decision-making agents in games (Silver et al., 2016; Berner et al., 2019; Vinyals et al., 2019; Mnih, 2016), robotics (Kalashnikov et al., 2018; Akkaya et al., 2019; Kumar et al., 2021; Hwangbo et al., 2019), and more recently, in developing reasoning capabilities in large language models (LLMs) (Ouyang et al., 2022; Guo et al., 2025; Team et al., 2023). Two widely used algorithms in this category are Trust Region Policy Optimization (TRPO) (Schulman et al., 2015) and Proximal Policy Optimization (PPO) (Schulman et al., 2017), both designed to stabilize learning by constraining the magnitude of policy updates.
@@
Another related line of work is Evolutionary Reinforcement Learning (EvoRL) (Moriarty et al., 1999), which integrates population-based search with policy optimization to explore a broad range of candidate solutions. Historically, evolutionary computation (EC) has been employed to optimize neural network weights (Tang et al., 2020; Choromanski et al., 2018, 2019), architectures (Gaier & Ha, 2019), and hyperparameters (Jaderberg et al., 2017; Snoek et al., 2012). For instance, OpenAI ES (Chrabaszcz et al., 2018) applied a streamlined evolution strategy to Atari games. More recently, researchers have integrated gradient-based updates alongside evolutionary operations (e.g., mutation and crossover), improving scalability and leveraging the representational power of deep neural networks (Pourchot & Sigaud, 2019; Khadka et al., 2019). For example, (Khadka & Tumer, 2018) extends Deep Deterministic Policy Gradient (DDPG) by introducing evolutionary operations on a population of policies, combining the exploratory benefits of evolution with gradient-based optimization for more robust learning. However, these methods still struggle with efficiently maintaining and updating large agent populations within GPU constraints, leading to low-quality data generation and poor sample efficiency. In this work, we propose to combine the strengths of evolutionary search with on-policy updates, enabling higher asymptotic rewards while maintaining high sample efficiency.
Off-Policy Reinforcement Learning
Unlike on-policy algorithms, off-policy methods allow for policy updates using data collected from different or older policies, enabling more efficient reuse of past experiences. Classic approaches such as Q-learning and its deep variant Deep Q-Network (DQN) (Mnih, 2013) rely on a replay buffer to store state-action transitions for more stable learning. More recent continuous control algorithms, such as Deep Deterministic Policy Gradient (DDPG) (Lillicrap, 2015; Silver et al., 2014), Twin Delayed DDPG (TD3) (Fujimoto et al., 2018), and Soft Actor-Critic (SAC) (Haarnoja et al., 2018), extend these ideas within an actor-critic framework, further improving stability and performance. While off-policy methods are typically more sample-efficient than on-policy methods, they often suffer from distribution shift and overestimation of action values, which can hinder stable learning. To effectively amortize experience across all policies in a population while ensuring stable updates, we adopt the Split and Aggregate Policy Gradient (SAPG) formulation (Singla et al., 2024), which performs off-policy updates via importance sampling, enabling the aggregation of experience from individual follower policies into the on-policy data of a master policy.
@@ 3 Preliminaries
$L_{on}(\pi_{\theta})=\mathop{\mathbb{E}}\limits_{\pi_{\text{old}}}\left[\min\left((r_{t}(\pi_{\theta}),\,\text{clip}\left(r_{t}(\pi_{\theta}),1-\epsilon,1+\epsilon\right)\right)A^{\pi_{old}}_{t}\right],$ (2)
@@
is one of the most well-known Evolutionary Algorithms (EAs). It relies on three key operators: selection, mutation, and crossover. A core concept in genetic algorithms is the population, where each individual represents a potential solution evaluated using a fitness function. The classic genetic algorithm emphasizes crossover as the primary mechanism for exploration (Hoffmeister & Bäck, 1990) and is usually applied to deal with the problem of hyper-parameter tuning in reinforcement learning.
@@ 4 Approach
EPO (As shown in Figure 1) begins by initializing a population of agents, where each agent is represented by a unique latent embedding (gene), while all agents share a common actor-critic network (gene-expression rules). Among these agents, one is designated as the “master” agent. Each agent then interacts with the environment for a full episode, and its fitness is determined by the cumulative reward obtained during that episode. A selection operator then determines which agents survive to the next generation, with survival probability proportional to their relative fitness scores. The embeddings of the surviving agents undergo mutation and crossover to form the next generation, while a small subset of top-performing agents (“elites”) is preserved unchanged to maintain stability.
To further enhance learning efficiency, we adopted a hybrid policy update strategy (Singla et al., 2024). The master agent updates its parameters using both (i) on-policy data collected in the most recent episode and (ii) off-policy data sampled from the trajectories of the entire population. Meanwhile, each non-master agent is updated using only its own on-policy data. This hybrid learning process ensures both efficient exploration and sample reuse, making EPO well-suited for large-scale reinforcement learning. In the following sections, we describe each component of EPO in detail.
@@ 4.1 Genetic Algorithm for Agent Selection
The design philosophy of our algorithm is twofold: (1) to generate diverse yet bounded data for training the master agent and (2) to ensure that non-master agents are updated efficiently, enabling them to produce useful yet diverse data. Drawing inspiration from biological evolution, we propose initializing a population of agents ($\{\pi_{k}\}_{k=1}^{K}$), where each agent is associated with a unique latent embedding ($\{\phi_{k}\}_{k=1}^{K}$). All agents share a common actor-critic network, parameterized by $\theta$ and $\psi$, respectively. The actor is conditioned on both the current state and its latent embedding to generate actions, while the value function is similarly conditioned to estimate value functions. Prior works have also leveraged latent-conditioned policies to encourage diverse behaviors (Ebert et al., 2021; Eysenbach et al., 2018). Among these agents, one is designated as the master agent—for simplicity, we define the master agent as $\pi_{1}(\theta,\psi,\phi_{1})$, though the general framework allows for alternative assignments.
To update non-master agents, we employ a genetic algorithm (GA) that modifies only the latent embeddings while keeping network parameters fixed. Each agent interacts with the environment for a full episode, and its fitness is determined by the cumulative reward obtained during that episode. A selection operator determines which agents survive to the next generation, with survival probability proportional to their relative fitness scores (i.e., cumulative reward). Specifically, we retain the top $x$ ranked non-master agents ($x\in[2,K]$) as elites, forming a set $X$, which is preserved unchanged and carried over to the next generation set $Y$. We then randomly select an agent from both $X$ and $Y$ to undergo crossover and mutation. While multiple crossover methods exist (e.g., k-point crossover, uniform crossover), we use a simple averaging method for efficiency. Given two selected latent embeddings, $\phi_{i}\in X$ and $\phi_{j}\in Y$, the crossover result is computed as: $\phi^{\prime}=1/2\times(\phi_{i}+\phi_{j})$. Following crossover, a mutation step is applied by adding Gaussian noise: $\phi^{\prime\prime}=\phi^{\prime}+\epsilon,\epsilon\sim\mathcal{N}(0,\sigma)$. The mutated latent embedding $\phi^{\prime\prime}$ is then added to $Y$ for the next generation.
@@ 4.2 Hybrid-Policy Optimization
Different agents generate diverse data through varied behaviors, which can be beneficial if the master agent can effectively learn from all available data. However, since data distributions collected by different agents may vary significantly, incorporating off-policy data requires careful correction. Following (Meng et al., 2023; Singla et al., 2024), we apply importance sampling to reweight updates from different policies. Specifically, we sample $|S_{1}|$ transitions from $\cup_{k=2}^{K}S_{k}$ collected by agents $\{\pi_{k}\}_{k=2}^{K}$ to get $S^{\prime}_{1}$. To update the master agent $\pi_{1}$ using $S^{\prime}_{1}$, we define the off-policy loss as:
$L_{\text{off}}(\pi_{1},S^{\prime}_{1})=\frac{1}{|S^{\prime}_{1}|}\mathop{\mathbb{E}}\limits_{(s,a)\sim S^{\prime}_{1}}[\min(r_{\pi_{1}}(s,a),$ (4)
$\quad\operatorname{clip}(r_{\pi_{1}}(s,a),\mu(1-\epsilon),\mu(1+\epsilon)))A^{\pi_{1},\text{old}}(s,a)]$
@@
13:        Select top $x$ as elites to form Set $X$, append to $Y$
@@
15:           Crossover between a randomly sampled $\phi_{i}\in X$ and $\phi_{j}\in Y$ to generate $\phi^{\prime}$
16:           Mutate($\phi^{\prime}$) and append it to $Y$
17:        end while
18:     end if
19:     for k = 1, …, $K$ do
20:        $R_{k}\leftarrow CollectData(E^{k},\theta,\psi,\phi_{k})$
21:     end for
22:     $L\leftarrow 0$
23:     Sample $|S_{1}|$ transitions from $\cup_{k=2}^{K}S_{k}$ to get $S_{1}^{\prime}$
24:     $L\leftarrow L+$ OffPolicyLoss($S_{1}^{\prime}$)
25:     for k = 1, …, $K$ do
26:        $L\leftarrow L+$ OnPolicyLoss($S_{k}$)
27:     end for
28:     Update $\theta$, $\psi$, $\{\phi_{k}\}_{k=1}^{K}$
@@
30:  Return $\theta$, $\psi$, $\phi_{1}$

### v2

@@ Abstract
On-policy reinforcement learning (RL) algorithms are widely used for their strong asymptotic performance and training stability, but they struggle to scale with larger batch sizes, as additional parallel environments yield redundant data due to limited policy-induced diversity. In contrast, Evolutionary Algorithms (EAs) scale naturally and encourage exploration via randomized population-based search, but are often sample-inefficient. We propose Evolutionary Policy Optimization (EPO), a hybrid algorithm that combines the scalability and diversity of EAs with the performance and stability of policy gradients. EPO maintains a population of agents conditioned on latent variables, shares actor-critic network parameters for coherence and memory efficiency, and aggregates diverse experiences into a master agent. Across tasks in dexterous manipulation, legged locomotion, and classic control, EPO outperforms state-of-the-art baselines in sample efficiency, asymptotic performance, and scalability. For visualizations of the learned policies, please visit: https://yifansu1301.github.io/EPO/.
@@ 1 Introduction
Reinforcement learning (RL) has become a powerful paradigm for training autonomous decision-making agents across a wide range of domains, including games [40, 2, 47, 30], robotics [19, 1, 22, 17], and large-language-model alignment [33, 13, 46]. Like all trial-and-error methods, RL is inherently sample-inefficient, and the problem is particularly acute for on-policy algorithms, which discard past experience after each update.
Despite this drawback, on-policy model-free methods—commonly referred to as policy gradients [38, 30]—are widely adopted in real-world applications [40, 2, 19, 1, 33]. Their popularity stems from their ability to achieve higher asymptotic returns and stable performance, making them the method of choice in data-rich settings such as simulations and game engines [26], where large batches are easy to generate. Unfortunately, recent works [41, 34] show that policy-gradient algorithms do not scale well with larger batch sizes: because data are collected from the current policy, adding more parallel environments does not guarantee greater diversity. Instead, the data distribution quickly converges, so simply running more environments on additional GPUs yields highly correlated trajectories rather than useful variety.
Evolutionary algorithms (EAs) contribute the complementary strength of explicit diversity through random perturbations, making them natural partners for RL. Their combination—evolutionary RL (EvoRL) [31, 36]—maintains a population of policies and refines them with evolutionary operators. Yet EvoRL has not been widely adopted. Classic EvoRL relies on gradient-free evolutionary strategies (ES) [20, 3, 48], whose simplest variants reduce to random search [27]: they parallelize well but remain extremely sample-inefficient. Methods such as CEM-RL [35] introduce off-policy updates to improve efficiency but degrade asymptotic performance owing to distribution shift. Population-Based Training (PBT) [34, 18] ensembles multiple on-policy agents, yet each policy evolves in isolation, wasting the opportunity to amortize experience across the population.
@@
Can we design an algorithm that (i) scales with the number of parallel environments, (ii) learns efficiently from shared experience, and (iii) preserves the high asymptotic returns and stability of on-policy methods? We answer in the affirmative. We introduce Evolutionary Policy Optimization (EPO), a novel policy-gradient algorithm that merges the strengths of EA and policy gradients: EA supplies diverse experience during massively parallel training, and experience from all agents is amortized to train a master agent via the split-and-aggregate policy-gradient (SAPG) formulation [41], which performs off-policy updates with importance sampling to fold follower data into the master’s on-policy set.
@@
EPO maintains a population of agents alongside a master agent, all sharing the same network weights (analogous to gene-expression rules). This design prevents parameter bloat as the population scales, enabling efficient training and storage of large networks. Each agent is conditioned on a unique latent variable (“gene”), ensuring behavioral diversity while preserving coherence. Periodically, Darwinian selection [8] is applied: low performers are removed, and elite agents are retained. These elites undergo crossover and mutation, injecting controlled diversity without excessive divergence. All agents are updated synchronously with policy gradients using shared reward signals. In every iteration, EPO aggregates experience from all agents into the master via SAPG updates [41], enabling the master to learn from diverse, high-quality data and fully exploit massive parallelism.
@@
We evaluate EPO on several challenging and realistic environments spanning manipulation (Isaac Gym Allegro-Kuka [26]), locomotion (ANYmal Walking [16], Unitree A1 Parkour [4, 49]), and classic control benchmarks (DeepMind Control Suite [45]). Across this suite, EPO significantly outperforms prior state-of-the-art RL methods. Moreover, our scaling-law analysis shows that EPO effectively scales with increased computational resources, making it a promising approach for large-scale RL training.
@@ 2 Related Work
On-policy approaches [43] update the policy using data collected by the current policy, ensuring that the training data remains closely aligned with the policy’s behavior. This strategy often leads to higher asymptotic rewards and stable performance, but at the cost of sample inefficiency, as past experiences from older policies are discarded. Despite this drawback, on-policy methods are the preferred choice in settings where data is abundant, such as simulations and game engines, where large training batches can be efficiently generated. In practice, on-policy RL has proven highly effective for training autonomous decision-making agents in games [40, 2, 47, 30], robotics [19, 1, 22, 17], and more recently, in developing reasoning capabilities in large language models (LLMs) [33, 13, 46]. Two widely used algorithms in this category are Trust Region Policy Optimization (TRPO) [37] and Proximal Policy Optimization (PPO) [38], both designed to stabilize learning by constraining the magnitude of policy updates.
Off-Policy Reinforcement Learning
Unlike on-policy algorithms, off-policy methods allow for policy updates using data collected from different or older policies, enabling more efficient reuse of past experiences. Classic approaches such as Q-learning and its deep variant Deep Q-Network (DQN) [29] rely on a replay buffer to store state-action transitions for more stable learning. More recent continuous control algorithms, such as Deep Deterministic Policy Gradient (DDPG) [25, 39], Twin Delayed DDPG (TD3) [11], and Soft Actor-Critic (SAC) [14], extend these ideas within an actor-critic framework, further improving stability and performance. While off-policy methods are typically more sample-efficient than on-policy methods, they often suffer from distribution shift and overestimation of action values, which can hinder stable learning. To effectively amortize experience across all policies in a population while ensuring stable updates, we adopt the Split and Aggregate Policy Gradient (SAPG) formulation [41], which performs off-policy updates via importance sampling, enabling the aggregation of experience from individual follower policies into the on-policy data of a master policy.
@@
Another related line of work is Evolutionary Reinforcement Learning (EvoRL) [31], which integrates population-based search with policy optimization to explore a broad range of candidate solutions. Historically, evolutionary computation (EC) has been employed to optimize neural network weights [44, 5, 6], architectures [12], and hyperparameters [18, 42]. For instance, OpenAI ES [7] applied a streamlined evolution strategy to Atari games. More recently, researchers have integrated gradient-based updates alongside evolutionary operations (e.g., mutation and crossover), improving scalability and leveraging the representational power of deep neural networks [35, 21]. For example, [20] extends Deep Deterministic Policy Gradient (DDPG) by introducing evolutionary operations on a population of policies, combining the exploratory benefits of evolution with gradient-based optimization for more robust learning. However, these methods remain vulnerable to distribution shift, which undermines their asymptotic performance. In this work, we propose to combine the strengths of evolutionary search with on-policy updates, enabling higher asymptotic rewards while maintaining high sample efficiency.
@@ 3 Preliminaries
$L_{on}(\pi_{\theta})=\mathop{\mathbb{E}}\limits_{\pi_{\text{old}}}\left[\min\left(r_{t}(\pi_{\theta})A^{\pi_{old}}_{t},\,\text{clip}\left(r_{t}(\pi_{\theta}),1-\epsilon,1+\epsilon\right)A^{\pi_{old}}_{t}\right)\right],$ (2)
@@
is one of the most well-known Evolutionary Algorithms (EAs). It relies on three key operators: selection, mutation, and crossover. A core concept in genetic algorithms is the population, where each individual represents a potential solution evaluated using a fitness function. The classic genetic algorithm emphasizes crossover as the primary mechanism for exploration [15] and is usually applied to deal with the problem of hyper-parameter tuning in reinforcement learning.
@@ 4 Approach
EPO (As shown in Figure 1) begins by initializing a population of agents, where each agent is represented by a unique latent embedding (gene), while all agents share a common actor-critic network (gene-expression rules). Among these agents, one is designated as the “master” agent. Each agent then interacts with the environment for a full episode, and its fitness value (defined differently for each task) is obtained during that episode. A selection operator then determines which agents survive to the next generation based on their fitness scores. The embeddings of the surviving agents undergo mutation and crossover to form the next generation, while a small subset of top-performing agents (“elites”) is preserved unchanged to maintain stability.
To further enhance learning efficiency, we adopted a hybrid policy update strategy [41]. The master agent updates its parameters using both (i) on-policy data collected in the most recent episode and (ii) off-policy data sampled from the trajectories of the entire population. Meanwhile, each non-master agent is updated using only its own on-policy data. This hybrid learning process ensures both efficient exploration and sample reuse, making EPO well-suited for large-scale reinforcement learning. In the following sections, we describe each component of EPO in detail.
@@ 4.1 Genetic Algorithm for Agent Selection
The design philosophy of our algorithm is twofold: (1) to generate diverse yet bounded data for training the master agent and (2) to ensure that non-master agents are updated efficiently, enabling them to produce useful yet diverse data. Drawing inspiration from biological evolution, we propose initializing a population of agents ($\{\pi_{k}\}_{k=1}^{K}$), where each agent is associated with a unique latent embedding ($\{\phi_{k}\}_{k=1}^{K}$). All agents share a common actor-critic network, parameterized by $\theta$ and $\psi$, respectively. The actor is conditioned on both the current state and its latent embedding to generate actions, while the value function is similarly conditioned to estimate value functions. Prior works have also leveraged latent-conditioned policies to encourage diverse behaviors [9, 10]. Among these agents, one is designated as the master agent—for simplicity, we define the master agent as $\pi_{1}(\theta,\psi,\phi_{1})$, though the general framework allows for alternative assignments.
To update non-master agents, we employ a genetic algorithm (GA) that modifies only the latent embeddings while keeping network parameters fixed. Each agent interacts with the environment for a full episode, and gets its fitness values. A selection operator determines which agents survive to the next generation, with survival probability proportional to their relative fitness scores (i.e., cumulative reward, success rate, etc.). Specifically, we retain the top $x$ ranked non-master agents ($x\in[2,K]$) as elites, forming a set $X$, which is preserved unchanged and carried over to the next generation set $Y$. We then randomly select two agents from $X$ to undergo crossover and mutation. While multiple crossover methods exist (e.g., k-point crossover, uniform crossover), we use a simple averaging method for efficiency. Given two selected latent embeddings, $\phi_{i},\phi_{j}$, where $\pi_{i}(\theta,\psi,\phi_{i}),\pi_{j}(\theta,\psi,\phi_{j})\in X$, the crossover result is computed as: $\phi^{\prime}=1/2\times(\phi_{i}+\phi_{j})$. Following crossover, a mutation step is applied by adding Gaussian noise: $\phi^{\prime\prime}=\phi^{\prime}+\epsilon,\epsilon\sim\mathcal{N}(0,\sigma^{2})$. The new agent $\pi(\theta,\psi,\phi^{\prime\prime})$ is then added to $Y$ for the next generation.
@@
, where $f_{k}$ represent the fitness score of agent $\pi_{k}$. A simpler alternative is to apply selection, crossover, and mutation at fixed intervals—e.g., every set number of environment steps.
@@ 4.2 Hybrid-Policy Optimization
Different agents generate diverse data through varied behaviors, which can be beneficial if the master agent can effectively learn from all available data. However, since data distributions collected by different agents may vary significantly, incorporating off-policy data requires careful correction. Following [28, 41], we apply importance sampling to reweight updates from different policies. Specifically, we sample $|S_{1}|$ transitions from $\cup_{k=2}^{K}S_{k}$ collected by agents $\{\pi_{k}\}_{k=2}^{K}$ to get $S^{\prime}_{1}$. To update the master agent $\pi_{1}$ using $S^{\prime}_{1}$, we define the off-policy loss as:
$L_{\text{off}}(\pi_{1},S^{\prime}_{1})=\frac{1}{|S^{\prime}_{1}|}\mathop{\mathbb{E}}\limits_{(s,a)\sim S^{\prime}_{1}}[\min(r_{\pi_{1}}(s,a)A^{\pi_{1},\text{old}}(s,a),\quad\operatorname{clip}(r_{\pi_{1}}(s,a),\mu(1-\epsilon),\mu(1+\epsilon))A^{\pi_{1},\text{old}}(s,a))]$ (4)
@@
13:        Select top $x$ agents as elites to form Set $X$, append to $Y$
@@
15:           ϕsuperscriptitalic-$\phi^{\prime}\leftarrow$Crossover between randomly sampled $\phi_{i},\phi_{j}$,where $\pi_{i}(\theta,\psi,\phi_{i}),\pi_{j}(\theta,\psi,\phi_{j})\in X$
16:           ϕ′′superscriptitalic-$\phi^{\prime\prime}\leftarrow$ Mutate($\phi^{\prime}$)
17:           Append $\pi(\theta,\psi,\phi^{\prime\prime})$ to $Y$
18:        end while
19:     end if
20:     $\{\pi_{i}\}_{i=2}^{K}\leftarrow Y$
21:     for k = 1, …, $K$ do
22:        CollectData$(E^{k},\pi_{k})$
23:     end for
24:     $L\leftarrow 0$
25:     Sample $|S_{1}|$ transitions from $\cup_{k=2}^{K}S_{k}$ to get $S_{1}^{\prime}$
26:     $L\leftarrow L+$ OffPolicyLoss($S_{1}^{\prime}$)
27:     for k = 1, …, $K$ do
28:        $L\leftarrow L+$ OnPolicyLoss($S_{k}$)
@@
30:     Update $\theta$, $\psi$, $\{\phi_{k}\}_{k=1}^{K}$
31:  end for
32:  Return $\theta$, $\psi$, $\phi_{1}$